what is the factor of 16m-28n?
what is the factor of 63v - 7w + 28?

Answer by checkley71(8403)

16M-28N=4(4M-7N) 4 IS THE ONLY FACTOR.
63V-7W+28=7(9V-W+4) 7 IS THE ONLY FACTOR.